OpenStudy (anonymous):

I'm not sure if I got this right -> Present the following into an algerbriac expression: A number doubles every time another number increases by one. Let x be the first number and y the second. What's the value of x if y increases by 16, assuming x = 2 and y = 1 wherein x = 2y?

OpenStudy (anonymous):

I got the following: \[x = 2(y+n)\]\[x = 2(1+16) = 2(17) = 34\]

OpenStudy (anonymous):

thats correct! :)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

no prb